OpenStudy (anonymous):

How do I find the inverse of this relation: y=3(2)^x???

OpenStudy (anonymous):

\[y=3*2^x\]

OpenStudy (anonymous):

this is the question ?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

ok divide both sides by 3 (y/3) = 2^x now introduce ln in both sides.. does it help?

OpenStudy (anonymous):

This is what I've done so far: y=3(2)^x x = 3(2)^y x/3 = 2^y logx/3 = log2^y logx/3 = ylog2 log(x/3)/log2 = y Using change of base formula: log[2](x/3).... but it's not right :(

OpenStudy (anonymous):

the correct answer states, \[y = \sqrt[3]{2\div3}\]

OpenStudy (anonymous):

i mean, \[y = \sqrt[3]{2}\div3 \]

OpenStudy (anonymous):

it is the right answer .. something wrong with the answer that you say its correct

OpenStudy (anonymous):

really?.. oh, okay. thank you!

OpenStudy (anonymous):

yw